OpenStudy (anonymous):

please help me with this !! find the range of (x^2+3x+10)/(x^2+6x+5) Ans: (-infinity, -3/4-sqrt(5/2)] U [-3/4+sqrt(5/2),infinity)

OpenStudy (anonymous):

\[y= \frac{x^2+3x+10}{x^2+6x+5}\]

OpenStudy (anonymous):

this is going to be a real pain i think you have to solve this for \(x\) and see what you get in terms of \(y\)

OpenStudy (campbell_st):

well you can factor the numerator and denominator... then look at the restictions in the domain... then see whats happening to the range close to those values.

OpenStudy (anonymous):

@satellite73 i did solve but not getting the right answer !!

OpenStudy (anonymous):

This isn't your typical well known functions like y = x^2, or y = sin(x), etc, that we can easily find the range with little to no calculation. So your only best bet is to rely on technology

OpenStudy (anonymous):

Type that function in wolfram alpha and you'l have the answer.

OpenStudy (anonymous):

@sourwing but i need to work it out !! ..Tried wolfram long back

OpenStudy (campbell_st):

why not just graph it... and see whats happening...

OpenStudy (anonymous):

You'll need calculus to find the range, which requires you to find increasing/decreasing, concave up/concave down, vertical/horizontal asymptotes, etc,.. You'll probably have to spend like an hour if you do it by hand.

OpenStudy (campbell_st):

here's the graph

OpenStudy (anonymous):

see what i mean? technology will give you answer right away. btw, horizontal asymptote is y = 1 :DD